A Triangle has a base that is six times the length of its height.
If the area of the triangle is 192 cm^2, find the height and length of the base (showing your reasoning)

Answer by mananth(16946)

A Triangle has a base that is six times the length of its height.
If the area of the triangle is 192 cm^2, find the height and length of the base
area of triangle = 1/2 * b *h
b= 6h
192= 1/2 * h * 6h
192=3h^2
/3
192/3=h^2
64=h^2
h= +/-8
ignore negative
h= 8 cm
and base = 6h = 48cm